OpenStudy (anonymous):

my sister is in 7th grade..needs help with this..please explain to her...thank u!!! the legs of a right triangle measure 24 feet and 45 feet. what is the length of the hypotenuse

Mehek (mehek14):

Does she know the Pythagorean Theorem?

Mehek (mehek14):

\(\text{a^2 + b^2 = c^2}\)

OpenStudy (anonymous):

she said yes

OpenStudy (anonymous):

@Mehek14

Mehek (mehek14):

ok so tell her to find what \(\text{24^2 + 45^2}\) = ?